[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Oct 24, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Oct 24 17:44:14 UTC 2017


HSU Shift Log for Tue Oct 24 2017 13:44:07 GMT-0400 (EDT)

Shift: H

Forecaster: R Pasch

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: None

AWIPS-2/NCP issues: None

Other Techincal issues: 00Z ECMWF/ECENS arrived 3.5 hours late

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171024/473734af/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list